A 9.00 L tank at 0.3 degrees C is filled with 12.8 g if chlorine pentafluoride gas and 17.1 g of sulfur tetrafluoride gas. You can assume both gases behave as ideal gases under these conditions.

Calculate the mole fraction of each gas. Round each of your answers to 3 significant digits.


chlorine pentafluoride =

sulfur tetrafluoride =


1
Expert's answer
2021-03-15T09:46:22-0400

M(chlorine pentafluoride) = 130.44 g/mol

n(chlorine pentafluoride) "= \frac{12.8}{130.44} = 0.098 \;mol"

M(sulfur tetrafluoride) = 108.07 g/mol

n(sulfur tetrafluoride) "= \frac{17.1}{108.07} = 0.158 \;mol"

Mole fraction of chlorine pentafluoride "= \frac{0.098}{0.158+0.098} = 0.382"

Mole fraction of sulfur tetrafluoride "= \frac{0.158}{0.158+0.098} = 0.617"